# SegMASt3R: Geometry Grounded Segment Matching

### v1

@@ 1 Introduction
Segment matching establishes correspondences between coherent regions—objects, parts, or semantic segments—across images. It underpins video object tracking [16], scene-graph construction [17, 22], and instance-level SLAM [25, 43]. Because it matches extended structures rather than sparse, texture-sensitive points, it is more robust to noise, occlusion, and appearance change. Mapping structured regions instead of isolated pixels also boosts interpretability and allows geometric or semantic priors to be injected, enabling higher-level reasoning about scene content.
Segment matching degrades sharply under wide‐baseline conditions, where images of the same scene are taken from widely separated viewpoints that introduce severe perspective, scale, and $\sim\!180^{\circ}$ rotation changes [33]. Such cases arise in long-term video correspondence and robotic navigation, and demand models that reason over global 3D structure and enforce geometric consistency. Current approaches that depend on features from pre-trained encoders like DINOv2 [29] or ViT [10] often mis-match repetitive patterns or fail to link drastically different views of the same object.
@@
We adapt MASt3R for segment matching by appending a lightweight segment-feature head that transforms patch-level embeddings into segment-level descriptors. Given an image pair, these descriptors are matched to establish segment correspondences. The head is trained with a contrastive objective patterned after SuperGlue [34]. Experiments show that our approach surpasses strong baselines—including SAM2’s video propagator, which is trained on far larger datasets—and state-of-the-art local-feature methods. Finally, we demonstrate its practical utility in two downstream tasks: object-relative navigation and instance-level mapping.
@@
  • We validate the practical utility of our approach by applying it to a real-world robotic navigation tasks, showcasing its effectiveness in downstream applications such as 3D Instance Mapping and Topological Navigation. Our methods outperforms competitors by significant margins showcasing the efficacy of our design choices
@@ 2 Related Work
Robust segment–level association has emerged as a crucial intermediate step for high–level vision tasks such as scene–graph construction, long–term object tracking, and multi–view instance association. While related sub-problems like video instance segmentation and object tracking have been studied extensively in recent works [45, 44, 40, 11, 30, 28, 23, 6], the broader challenge of matching segments across arbitrary viewpoints, modalities, and time remains comparatively under-explored. Large-scale class-agnostic segmentation models have begun to close this gap. The Segment Anything Model (SAM) [19] and its successor SAM2 [33] deliver high-quality masks and include a built-in propagator for associating masks across video frames. However, this propagation module is optimized for short temporal windows and does not explicitly enforce geometric consistency under wide baselines or substantial appearance changes.
@@
Recent methods address segment matching more directly. MASA [24] augments SAM’s rich object proposals with synthetic geometric transformations to learn instance correspondences, and DMESA [48] extends these ideas to dense matching with improved efficiency. Despite such progress, these approaches are still limited by 2D supervision. An alternative line of work predicts the degree of visual overlap between images [7, 13, 2]. By estimating shared content, these methods implicitly learn region correspondences, yet they also remain confined to 2D training signals.
@@
Sparse [9, 34, 39, 26, 35] and dense [12, 4] local feature matchers propagate pixel-level correspondences that can, in principle, transfer segment labels between views. Nonetheless, like the previous categories, they are trained exclusively on image data and struggle with extreme viewpoint changes.
@@
Figure 1: Overview Pipeline overview: a wide‑baseline image pair is processed by a frozen MASt3R backbone to extract patch‑level features; masks from a parallel segmentation module or ground truth masks and these patch-level features are aggregated by the segment-feature head (denoted Segment in figure) to form per‑segment descriptors, which are then matched across images via a differentiable optimal‑transport layer to produce the final matched segments.
@@ 3 Method
Figure˜1 refers to the overview of our method. It builds upon the MASt3R [21] architecture by introducing a Feat2Seg Adapter that maps the patch-level features output by the MASt3R decoder to get segment-level features, which are subsequently matched via a differentiable optimal transport and a row-wise argmax yielding the final segmentation.
@@ 3.1 MASt3R Preliminaries
Next, a pair of CroCo [41, 42] style intertwined transformer decoders jointly refines the two feature sets. By alternating self‑ and cross‑attention, the decoders exchange information to capture both the relative viewpoints and the global 3‑D structure of the scene. We show via an ablation study, these cross-view aware decoders provide a significant boost to the segment matching in Section˜5
@@
Assuming, $H,W$ denote the height and width of the input images, the output geometry‑aware patch-level features are of size $(H/16,W/16,768)$. The original architecture subsequently uses two prediction heads to output dense point-maps, as well as pixel-level features, they are not shown in Figure˜1 as they are not used in our pipeline. These portions of the pipeline remain frozen and their outputs are used as is.
@@ 3.2 Segment-Feature Head: Segment-Aligned Features
The MASt3R decoder produces patch–level embeddings $V_{1},V_{2}$ of shape $(H/16,W/16,768)$. In the original MASt3R [21], a feature head upsamples these tensors to the resolution of the input image. For segment matching we introduce another head to transform the patch level features to $M$ segment features. This introduced head is realized as an MLP that upsamples the patch-level features $(V_{1},V_{2})$ to image resolution, yielding feature maps of size $(H,W,24)$, $24$ being the feature dimension. We denote this Feature-to-Segment head as the segment-feature head. In addition to the patch-features, the segment-feature takes as input $M$ image resolution segment masks,—obtained either from an external segmenter such as SAM2 [33] or from ground-truth annotations. Both the masks and feature maps are flattened along the spatial dimensions yielding flattened tensors.
@@
The resulting segment embeddings for the two images are denoted $\mathbf{G}_{1}$ and $\mathbf{G}_{2}$ in Figure˜1 and are fed to the differentiable matching layer described in Section 3.3. We set $M=100$, which we found sufficient to cover the entire image while keeping computation affordable.
@@ 3.3 Differentiable Segment Matching Layer
As, the focus of this paper is segment matching under the wide-baseline setting, it become critical for the method to be able to mark segments in the first image that don’t have an analogue in the second image as unmatched, following [34], we incorporate a dustbin row and column to the affinity matrix $\mathbf{S}$ that absorbs non-matches. We augment (2) by concatenating an additional row and column initialized with a learnable logit $\alpha\in\mathbb{R}$,
@@ 4.1 Datasets
ScanNet++ [46]. ScanNet++ contains 1 006 indoor scenes captured with DSLR images and RGB-D iPhone streams, all co-registered to high-quality laser scans. The dataset supplies 3D semantic meshes, 2D instance masks, and accurate camera poses, which we utilize in our pipeline. We train on 860 k image pairs from 140 scenes (MASt3R split) and evaluate on 8 k pairs from 36 validation scenes, sampled with a fixed seed (42) and balanced across scenes and four pose bins: $0^{\circ}\!-\!45^{\circ}]$, $45^{\circ}\!-\!90^{\circ}]$, $90^{\circ}\!-\!135^{\circ}]$, and $135^{\circ}\!-\!180^{\circ}]$, defined by the rotational geodesic distance between camera orientations.
Replica [38]. The Replica dataset contain 18 high-quality, photo-realistic indoor room reconstructions in the form of dense meshes replete with per-primitive semantic class and instance information. In particular, we use a pre-rendered version of this dataset from Semantic-NeRF [49], which directly provides poses, RGB-D sequences, and per-frame semantic masks. For evaluation, we employ the same pose-binning strategy as described above and randomly sample 3200 image pairs across 8 scenes, again ensuring uniform sampling across both scenes and pose-bins.
MapFree Visual Re-localization Dataset [2]. MapFree Dataset is a highly challenging benchmark for metric-relative camera pose estimation. It comprises 655 diverse outdoor scenes—sculptures, fountains, murals, and other points of interest—designed to challenge models to severe stress tests, including extreme viewpoint changes, varying visual conditions between query and reference images, and geometric ambiguities such as symmetric structures. We limit our use of this dataset to just qualitative evaluation, as it does not provide 3D instance segmentation ground truth.
@@ 4.2 Baselines
Local Feature Matching (LFM). Although few models target segment matching directly, a wide range of local feature matchers (LFMs) exists at different densities—sparse SuperPoint [9], semi-dense LoFTR [39], and dense RoMa and MASt3R [12, 21]. We harness these state-of-the-art LFMs via the EarthMatch toolkit [5] to obtain segment correspondences. Each matched keypoint pair votes for the source and target masks that contain its coordinates, populating a vote matrix of size $M\times N$ (with $M$ and $N$ segments in the two images). Correspondences are taken as the highest-scoring entries of this matrix; the full algorithm is provided in the supplementary.
Segment Matching (SegMatch). We also compute segment matches from dense features of two strong vision encoders, DINOv2 [29] and MASt3R [21]. A segment descriptor is the masked average of its feature map at a suitable resolution; cosine similarity between all source–target descriptors yields a match matrix, from which one-to-one links are selected via row-wise argmax. We downsample the masks to the encoder output resolution to extract features. We further benchmark SegVLAD [14], which aggregates features from neighbouring segments.
@@ 4.3 Evaluation Metrics
We report two complementary measures of segment correspondence quality - AUPRC and Recall. Area Under the Precision–Recall Curve (AUPRC) integrates precision over the entire recall axis, providing a threshold–independent summary that is particularly informative under the high class–imbalance characteristic of segment matching between image pairs. Recall@$k$ ($R@k$) denotes the fraction of query segments whose ground-truth counterpart is found within the top $k$ ranked candidates, thus gauging how effectively the method surfaces correct matches among its highest-confidence predictions. Additional dataset, baseline and experimental details can be found in the supplementary, along with more qualitative results.
@@ 5 Results
In Table 1, we compare our proposed method SegMASt3R with the state-of-the-art methods using two categories of approaches in the literature: the well-established local feature matching (LFM) based on sparse or dense keypoint correspondences, and the recently emerging open-set instance association based on segment matching (SegMatch). It can be observed that SegMASt3R outperforms all the baselines for all the pose bins with a huge margin. Amongst the LFM techniques, dense matchers (RoMa and MASt3R) outperform sparse matchers (SP-LG and GiM-DKM) on both the evaluation metrics for the task of segment matching. Notably, on the highly challenging wide-baseline settings, MASt3R outperforms other local matchers by a large margin. However, when the same backend features are aggregated at the segment-level, MASt3R’s performance deteriorates significantly, e.g., AUPRC drops from 52.9 to 41.4 on the 90-135 pose bin. This uncovers an inherent limitation: feature distinctiveness at the pixel level does not necessarily translate to the instance level. Considering the SegMatch techniques – which are not trained for the LFM task – it can be observed that DINOv2 and SegVLAD perform particulary well on R@5 metric, which aligns with their typical use for coarse retrieval [18, 14]. On the other hand, SAM2’s two-frame video propagation only works well for narrow-baseline matching, which can be expected as its training set is mostly comprised of dynamic object tracking. Overall, these results show that all the baseline methods lack on at least one of the fronts: pixel- vs segment-level distinctiveness, recall vs precision, and narrow- vs wide-baseline robustness. Our proposed method SegMASt3R achieves all these desirable properties with high performance across the board by learning segment-level representations with the training objective of instance association. In Section 5, we provide qualitative results which emphasize the ability of our method to address the problem of perceptual instance aliasing and instance matching under extreme viewpoint shifts.
@@
Furthermore, we test generalization to challenging outdoor scenes from the MapFree dataset [2], with results shown in Table˜3. Since MapFree lacks instance-level ground truth, we use SAM2’s video propagator on image sequences to generate a pseudo-ground truth for evaluation. Our indoor-trained model (SegMASt3R (SPP)) shows a regression in performance in comparison to DINOv2, highlighting the domain shift. However, this gap can be substantially closed by either re-training on MapFree data (SegMASt3R (MF)) or, even more simply by just recalibrating the single learnable dustbin parameter $\alpha$ using a simple grid-search over a small calibration set from the target domain (SegMASt3R (SPP, Dustbin MF)). This demonstrates the strong adaptability of our model’s learned geometric features.
@@
Method Train Set Overall IoU 0-45° 90-135° 135-180°
DINOv2 [29] Multiple 84.4 85.4 83.5 83.8
MASt3R (Vanilla) [21] Multiple 69.2 73.4 70.0 66.1
SegMASt3R (SPP) ScanNet++ 75.2 75.2 76.5 74.5
SegMASt3R (SPP, Dustbin MF) ScanNet++ 88.7 88.6 88.5 93.9
@@
ConceptGraph (MobileSAM Masks) [16] 11.84 / 28.43 20.31 / 43.79 8.63 / 22.82 8.07 / 22.83 9.46 / 24.73 12.23 / 34.34 5.83 / 12.96 7.83 / 23.82
ConceptGraph (GT Masks) [16] 43.53 / 69.68 22.48 / 40.71 43.46 / 60.69 32.06 / 53.44 39.63 / 68.22 44.89 / 69.64 17.96 / 36.53 25.93 / 43.63
@@
To decouple these factors and isolate the core matching performance, we adopt an evaluation protocol that does not penalize methods for AMG errors. Specifically, while all methods take noisy masks as input, the evaluation is performed only on the subset of ground-truth correspondences for which a valid match between predicted segments was possible. For a comprehensive analysis, we also report results for the evaluation which jointly assesses both matching and segmentation alignment in the supplementary material. As shown in Table 5, SegMASt3R maintains a substantial performance margin over all baselines across different evaluation thresholds. This result confirms that the strong geometric priors learned by our method provide significant robustness, enabling superior matching even when conditioned on inconsistent and noisy segment inputs.
@@
Recent works [16, 15] have explored the use of SAM2’s open-set, semantically-meaningful instance segmentation for topological mapping and navigation. These methods rely on accurate segment-level association. To show the benefits of our proposed method on complex downstream tasks such as navigation, we considered an object topology-based mapping and navigation method, RoboHop [15], and swapped its DINOv2-based segment matching with our proposed segment matcher. This segment matching is used by RoboHop’s localizer to match each of the object instances in the query image with those in the sub-map images, which is then used to estimate the current object’s distance to the goal and correspondingly obtain a control signal. We used the val set of the ImageInstanceNav [20] dataset, which comprises real-world indoor scenes from HM3Dv0.2 [32]. We followed [31] for creating map trajectories and evaluating navigation performance.
In Table 6, we report navigation success for (vanilla) RoboHop and an enhanced version of it that uses SegMASt3R for segment association for localization and navigation. We use two metrics: SPL (Success weighted by Path Length) [1] and SSPL (soft SPL) [8]. Furthermore, we considered four different evaluation settings using two parameters that define the submap used for localizing the query image segments: Submap Span ($S_{s}$), which defines the total number of images sampled from the map based on the distance from the robot’s current position, and Submap Density ($S_{\rho}$), which defines a subsampling factor to uniformly skip map images. These parameter configurations aid in testing narrow- and wide-baseline matching as well as the ability to avoid false positives. Table˜6 shows that SegMASt3R consistently outperforms vanilla RoboHop. In particular, we achieve an absolute improvement of $27\%$ SPL on one of the hardest parameter settings: $S_{s}=16,S_{\rho}=0.25$, where only $4$ submap images are sampled due to a low density value. This shows that even with a very sparse submap, it is possible to maintain high navigation success rate, thus avoiding the typical trade-off between compute time and accuracy for the localizer. In Figure 2, we present a qualitative comparison between vanilla RoboHop’s DINOv2-based segment matching and that based on SegMASt3R: (left) the mismatch between the wall (orange) and the vanity cabinet for DINOv2 leads to an incorrect rotation towards the right, consequently leading to navigation failure, whereas SegMASt3R’s accurate segment matching (right) guides the agent into the bathroom and finally to the goal object.
@@
Figure 2: Segment Matching-Guided Navigation. (left) In vanilla RoboHop with DINOv2 based segment matching, a wall segment (orange) gets mismatched with the vanity cabinet and misguides the agent to move towards its right, leading to a navigation failure. (right) SegMASt3R correctly recognizes the same cabinet as well as other segments and guides the robot into the bathroom, and eventually to the goal. Note that the query and submap images vary across both the methods, as we manually probed the point of failure for the baseline and the nearest agent state for ours.
Impact of the Feature Encoder.

### v2

@@ Abstract
{NoHyper} Equal Contribution. Corresponding Author. Now at Google DeepMind.
@@ 1 Introduction
Segment matching establishes correspondences between coherent regions—objects, parts, or semantic segments—across images. It underpins video object tracking [19], scene-graph construction [20, 25], robot navigation [18, 17, 34], and instance-level SLAM [28, 47]. Because it matches extended structures rather than sparse, texture-sensitive points, it is more robust to noise, occlusion, and appearance change. Mapping structured regions instead of isolated pixels also boosts interpretability and allows geometric or semantic priors to be injected, enabling higher-level reasoning about scene content.
Segment matching degrades sharply under wide‐baseline conditions, where images of the same scene are taken from widely separated viewpoints that introduce severe perspective, scale, and up to $180^{\circ}$ rotation changes [37]. Such cases arise in long-term video correspondence and robotic navigation, and demand models that reason over global 3D structure and enforce geometric consistency. Current approaches that depend on features from pre-trained encoders like DINOv2 [32] or ViT [12] often mismatch repetitive patterns or fail to link drastically different views of the same object.
@@
We adapt MASt3R for segment matching by appending a lightweight segment-feature head that transforms patch-level embeddings into segment-level descriptors. Given an image pair, these descriptors are matched to establish segment correspondences. The head is trained with a contrastive objective patterned after SuperGlue [38]. Experiments show that our approach surpasses strong baselines, including SAM2’s video propagator which is trained on far larger datasets and state-of-the-art local feature matching methods. Finally, we demonstrate its practical utility in two downstream tasks: object-relative navigation and instance-level mapping.
@@
  • We validate the practical utility of our approach by applying it to the downstream applications such as 3D instance mapping and object-relative topological navigation. Our method outperforms competitors by significant margins showcasing the efficacy of our design choices.
@@ 2 Related Work
Robust segment-level association has emerged as a crucial intermediate step for high-level vision tasks such as scene graph construction, long-term object tracking, and multi-view instance association. While related sub-problems like video instance segmentation and object tracking have been studied extensively in recent works [49, 48, 44, 13, 33, 31, 26, 7], the broader challenge of matching segments across arbitrary viewpoints, modalities, and time remains comparatively under-explored. Large-scale class-agnostic segmentation models have begun to close this gap. The Segment Anything Model (SAM) [22] and its successor SAM2 [37] deliver high-quality masks and include a built-in propagator for associating masks across video frames. However, this propagation module is optimized for short temporal windows and does not explicitly enforce geometric consistency under wide baselines or substantial appearance changes.
@@
Some recent methods address segment matching more directly. MASA [27] augments SAM’s rich object proposals with synthetic geometric transformations to learn instance correspondences, and DMESA [52] extends these ideas to dense matching with improved efficiency. Despite such progress, these approaches are still limited by 2D supervision. An alternative line of work predicts the degree of visual overlap between images [8, 15, 2]. By estimating shared content, these methods implicitly learn region correspondences, yet they also remain confined to 2D training signals.
@@
Sparse [11, 38, 43, 29, 39] and dense [14, 4] local feature matchers propagate pixel-level correspondences that can, in principle, transfer segment labels between views [17, 18]. Nonetheless, like the previous categories, they are trained exclusively on image data and struggle with extreme viewpoint changes.
@@
Figure 1: Pipeline Overview: An image pair is processed by a frozen MASt3R backbone to extract patch‑level features; segmentation masks are obtained either from a parallel segmentation module or ground truth annotation; the patch-level features are aggregated by the segment-feature head to form segment-level descriptors; and these descriptors are then matched across images via a differentiable optimal transport layer to produce segment-level correspondences.
@@ 3 Method
Figure˜1 provides an overview of our method. It builds upon the MASt3R [24] architecture by introducing a Feat2Seg Adapter that maps the patch-level features output by the MASt3R decoder to get segment-level features, which are subsequently matched via a differentiable optimal transport and a row-wise argmax yielding the final segment-level correspondences.
@@ 3.1 MASt3R Preliminaries
Next, a pair of CroCo [45, 46] style intertwined transformer decoders jointly refines the two feature sets. By alternating self‑ and cross‑attention, the decoders exchange information between inputs to capture both the relative viewpoints and the global 3‑D structure of the scene. We show via an ablation study, these cross-view aware decoders provide a significant boost to the segment matching in Section˜5
@@
Assuming, $H,W$ denote the height and width of the input images, the output geometry‑aware patch-level features are of size $(H/16,W/16,768)$. The original architecture subsequently uses two prediction heads to output dense point-maps, as well as pixel-level features, they are not shown in Figure˜1 as they are not used in our pipeline. These portions of the pipeline remain frozen and their outputs are used as is, we instead introduce a new segment-feature head.
@@ 3.2 Segment-Feature Head: Segment-Aligned Features
The MASt3R decoder produces patch–level embeddings $V_{1},V_{2}$ of shape $(H/16,W/16,768)$. In the original MASt3R [24], a feature head upsamples these tensors to the resolution of the input image. For segment matching we introduce another head to transform the patch level features to $M$ segment features. This introduced head is realized as an MLP that upsamples the patch-level features $(V_{1},V_{2})$ to image resolution, yielding feature maps of size $(H,W,24)$, $24$ being the feature dimension. We denote this Feature-to-Segment head as the segment-feature head. In addition to the patch-features, the segment-feature takes as input $M$ image resolution segment masks, obtained either from an external segmenter such as SAM2 [37] or from ground-truth annotations. Both the masks and feature maps are flattened along the spatial dimensions yielding flattened tensors.
@@
The resulting segment embeddings for the two images are denoted $\mathbf{G}_{1}$ and $\mathbf{G}_{2}$ in Figure˜1 and are fed to the differentiable matching layer described in Section 3.3. We set $M=100$ as an upper bound for batch processing. In practice, images typically contain 20-30 masks when training with ground truth annotations; we pad with zeros when fewer masks are present. At inference time, the number of masks can be set arbitrarily, independent of the training-time value of $M$.
@@ 3.3 Differentiable Segment Matching Layer
Following [38], we incorporate a dustbin row and column in the affinity matrix $\mathbf{S}$ to handle segments without correspondences in the other image, which is critical for wide-baseline matching. We augment (2) by concatenating an additional row and column initialized with a learnable logit $\alpha\in\mathbb{R}$, yielding $\tilde{\mathbf{S}}\in\mathbb{R}^{(M_{1}+1)\times(M_{2}+1)}$.
@@ 4.1 Datasets
ScanNet++ [50].
ScanNet++ contains 1 006 indoor scenes captured with DSLR images and RGB-D iPhone streams, all co-registered to high-quality laser scans. The dataset supplies 3D semantic meshes, 2D instance masks, and accurate camera poses, which we utilize in our pipeline. We train on 860 k image pairs from 140 scenes and evaluate on 8 k pairs from 36 validation scenes, sampled with a fixed seed (42) and balanced across scenes and four pose bins: $0^{\circ}\!-\!45^{\circ}]$, $45^{\circ}\!-\!90^{\circ}]$, $90^{\circ}\!-\!135^{\circ}]$, and $135^{\circ}\!-\!180^{\circ}]$, defined by the rotational geodesic distance between camera orientations.
Replica [42].
The Replica dataset contain 18 high-quality, photo-realistic indoor room reconstructions in the form of dense meshes replete with per-primitive semantic class and instance information. In particular, we use a pre-rendered version of this dataset from Semantic-NeRF [53], which directly provides poses, RGB-D sequences, and per-frame semantic masks. For evaluation, we employ the same pose-binning strategy as described above and randomly sample 3200 image pairs across 8 scenes, again ensuring uniform sampling across both scenes and pose-bins.
MapFree Visual Re-localization [2].
MapFree is a challenging benchmark for metric-relative pose estimation, featuring 655 diverse outdoor scenes (sculptures, fountains, murals) with extreme viewpoint changes, varying visual conditions, and geometric ambiguities. The training split contains 460 scenes with 0.5M images; we uniformly sample 50 scenes yielding 31K image pairs for training. Since test split ground truth is unavailable, we report results on the validation split using 7.8K pairs from 13 uniformly sampled scenes (out of 65 total).
@@ 4.2 Baselines
Local Feature Matching (LFM).
Although few models target segment matching directly, a wide range of local feature matchers (LFMs) exists at different densities: sparse SuperPoint [11], semi-dense LoFTR [43], and dense RoMa and MASt3R [14, 24]. We harness these state-of-the-art LFMs via the EarthMatch toolkit [5] to obtain segment correspondences. Each matched keypoint pair votes for the source and target masks that contain its coordinates, populating a vote matrix of size $M\times N$ (with $M$ and $N$ segments in the two images). Correspondences are taken as the highest-scoring entries of this matrix; the full algorithm is provided in the supplementary.
Segment Matching (SegMatch).
We also compute segment matches from dense features of two strong pre-trained vision encoders, DINOv2 [32] and MASt3R [24], as well as SAM2’s video propagator [37] to track masks across views. For feature based methods, masks are downsampled via nearest-neighbor interpolation to match feature resolution, and segment descriptors are computed via masked average pooling. Cosine similarity between descriptors yields a match matrix, from which one-to-one links are selected via mutual-check. We further benchmark against SegVLAD [16], which aggregates features from neighbouring segments for segment retrieval based visual place recognition.
@@ 4.3 Evaluation Metrics
We report two complementary measures of segment correspondence quality - AUPRC and Recall. Area Under the Precision–Recall Curve (AUPRC) integrates precision over the entire recall axis, providing a threshold–independent summary that is particularly informative under the high class–imbalance characteristic of segment matching between image pairs. Recall@$k$ ($R@k$) denotes the fraction of query segments whose ground-truth counterpart is found within the top $k$ ranked candidates, thus gauging how effectively the method surfaces correct matches among its highest-confidence predictions. Additional details for the dataset, baseline and experiments can be found in the supplementary, along with more qualitative results.
@@ 5 Results
In Table 1, we compare our proposed method SegMASt3R with the state-of-the-art methods using two categories of approaches in the literature: the well-established local feature matching (LFM) based on sparse or dense keypoint correspondences, and the recently emerging open-set instance association based on segment matching (SegMatch). It can be observed that SegMASt3R outperforms all the baselines for all the pose bins with a huge margin. Amongst the LFM techniques, dense matchers (RoMa and MASt3R) outperform sparse matchers (SP-LG and GiM-DKM) on both the evaluation metrics for the task of segment matching. Notably, on the highly challenging wide-baseline settings, MASt3R outperforms other local matchers by a large margin. However, when the same backend features are aggregated at the segment-level, MASt3R’s performance deteriorates significantly, e.g., AUPRC drops from 52.9 to 41.4 on the 90-135 pose bin. This uncovers an inherent limitation: feature distinctiveness at the pixel level does not necessarily translate to the instance level. Considering the SegMatch techniques which are not trained for the LFM task, it can be observed that DINOv2 and SegVLAD perform particulary well on R@5 metric, which aligns with their typical use for coarse retrieval [21, 16]. On the other hand, SAM2’s two-frame video propagation only works well for narrow-baseline matching, which can be expected as its training set is mostly comprised of dynamic object tracking. Overall, these results show that all the baseline methods lack on at least one of the fronts: pixel- vs segment-level distinctiveness, recall vs precision, and narrow- vs wide-baseline robustness. Our proposed method SegMASt3R achieves all these desirable properties with high performance across the board by learning segment-level representations with the training objective of instance association. In Section 5, we provide qualitative results which emphasize the ability of our method to address the problem of perceptual instance aliasing and instance matching under extreme viewpoint shifts.
@@
Furthermore, we test generalization to challenging outdoor scenes from the MapFree dataset [2], with results shown in Table˜3. Since MapFree lacks instance-level ground truth, we use SAM2’s video propagator on image sequences to generate a pseudo-ground truth for evaluation. Our indoor-trained model (SegMASt3R (SPP)) shows a regression in performance in comparison to DINOv2, highlighting the domain shift. However, this gap can be substantially closed by either re-training on MapFree data (SegMASt3R (MF)) or, even more simply by just recalibrating the single learnable dustbin parameter $\alpha$ using a grid-search over a small calibration set from the target domain (SegMASt3R (SPP, Dustbin MF)). This demonstrates the strong adaptability of our model’s learned geometric features.
@@
Method Train Set Overall IoU 0-45° 45-90° 90-135° 135-180°
DINOv2 [32] Multiple 84.4 85.4 85.2 83.5 83.8
MASt3R (Vanilla) [24] Multiple 69.2 73.4 69.8 70.0 66.1
SegMASt3R (SPP) ScanNet++ 75.2 75.2 74.6 76.5 74.5
SegMASt3R (SPP, Dustbin MF) ScanNet++ 88.7 88.6 88.6 88.5 88.9
@@
ConceptGraphs (MobileSAM Masks) [19] 11.84 / 28.43 20.31 / 43.79 8.63 / 22.82 8.07 / 22.83 9.46 / 24.73 12.23 / 34.34 5.83 / 12.96 7.83 / 23.82
ConceptGraphs (GT Masks) [19] 43.53 / 69.68 22.48 / 40.71 43.46 / 60.69 32.06 / 53.44 39.63 / 68.22 44.89 / 69.64 17.96 / 36.53 25.93 / 43.63
@@
To decouple these factors and isolate the core matching performance, we adopt an evaluation protocol that does not penalize methods for AMG errors. Specifically, while all methods take noisy masks as input, the evaluation is performed only on the subset of ground-truth correspondences for which a valid match between predicted segments was possible. As shown in Table 5, SegMASt3R maintains a substantial performance margin over all baselines across different evaluation thresholds. This result confirms that the strong geometric priors learned by our method provide significant robustness, enabling superior matching even when conditioned on inconsistent and noisy segment inputs.
@@
Recent works [19, 17, 18] have explored the use of SAM2’s open-set, semantically-meaningful instance segmentation for topological mapping and navigation. These methods rely on accurate segment-level association. To show the benefits of our proposed method on complex downstream tasks such as navigation, we considered an object topology-based mapping and navigation method, RoboHop [17], and swapped its SuperPoint+LightGlue based segment matching with our proposed segment matcher. This segment matching is used by RoboHop’s localizer to match each of the object instances in the query image with those in the sub-map images, which is then used to estimate the currently-viewed object’s path to the goal and correspondingly obtain a control signal. We used the val set of the ImageInstanceNav [23] dataset, which comprises real-world indoor scenes from HM3Dv0.2 [36]. We followed [34] for creating map trajectories and evaluating navigation performance.
In Table 6, we report navigation success for (vanilla) RoboHop and an enhanced version of it that uses SegMASt3R for segment association for localization and navigation. We use two metrics: SPL (Success weighted by Path Length) [1] and SSPL (soft SPL) [10]. Furthermore, we considered four different evaluation settings using two parameters that define the submap used for localizing the query image segments: Submap Span ($S_{s}$), which defines the total number of images sampled from the map based on the distance from the robot’s current position, and Submap Density ($S_{\rho}$), which defines a subsampling factor to uniformly skip map images. These parameter configurations aid in testing narrow- and wide-baseline matching as well as the ability to avoid false positives. Table˜6 shows that SegMASt3R consistently outperforms vanilla RoboHop. In particular, we achieve an absolute improvement of $27\%$ SPL on one of the hardest parameter settings: $S_{s}=16,S_{\rho}=0.25$, where only $4$ submap images are sampled due to a low density value. This shows that even with a very sparse submap, it is possible to maintain high navigation success rate, thus avoiding the typical trade-off between compute time and accuracy for the localizer. In Figure 2, we present a qualitative comparison between vanilla RoboHop’s segment matching and that based on SegMASt3R: (left) the mismatch between the wall (orange) and the vanity cabinet for the vanilla methods leads to an incorrect rotation towards the right, consequently leading to navigation failure, whereas SegMASt3R’s accurate segment matching (right) guides the agent into the bathroom and finally to the goal object.
@@
Figure 2: Segment Matching-Guided Navigation. (left) In vanilla RoboHop’s segment matching, a wall segment (orange) gets mismatched with the vanity cabinet and misguides the agent to move towards its right, leading to a navigation failure. (right) SegMASt3R correctly recognizes the same cabinet as well as other segments and guides the robot into the bathroom, and eventually to the goal. Note that the query and submap images vary across both the methods, as we manually probed the point of failure for the baseline and the nearest agent state for ours.
Impact of the Feature Encoder